You measure 40 textbooks' weights, and find they have a mean weight of 45 ounces. Assume the population standard deviation is 7.5 ounces. Based on this, construct a 95% confidence interval for the true population mean textbook weight.
Answer #1

Solution:

Confidence interval for Population mean is given as below:

Confidence interval = Xbar ± Z*σ/sqrt(n)

From given data, we have

Xbar = 45

σ = 7.5

n = 40

Confidence level = 95%

Critical Z value = 1.96

(by using z-table)

Confidence interval = Xbar ± Z*σ/sqrt(n)

Confidence interval = 45 ± 1.96*7.5/sqrt(40)

Confidence interval = 45 ± 2.3242

Lower limit = 45 - 2.3242 = 42.68

Upper limit = 45 + 2.3242 = 47.32

Confidence interval = (42.68, 47.32)

42.68 ounce < µ < 47.32 ounce